A rectangle is cut by 6 lines parallel and 8 lines perpendicular to the base. Find the total number of rectangles thus formed Answer by ikleyn(53751) 
 .
A rectangle is cut by 6 lines parallel and 8 lines perpendicular to the base.
Find the total number of rectangles thus formed
~~~~~~~~~~~~~~~~
First 6 cuts made 7 = 6+1 strips parallel to the base.
Next 8 cuts perpendicular to the base cut each of 7 parallel strips in 9 = 8+1 small rectangles.
Total number of small cut rectangles is the product 7*9 = 63. ANSWER